OpenStudy (anonymous):

another one, limit x->-pi, (sqrtx+4) cos(x+pi) help!!

OpenStudy (zarkon):

this? \[\lim_{x\to -\pi}\sqrt{x+4}\cdot \cos(x+\pi)\]